OpenStudy (anonymous):

Can someone please help me? Write the expression as either the sine, cosine, or tangent of a single angle. cos(pi/3)cos(pi/5)+sin(pi/3)sin(pi/5)

OpenStudy (anonymous):

Using cos(A-B) = cos A cos B + sin A sin B your A = pi/3 your B = pi/5 So this can be written as cos (pi/3 - pi/5) or just cos (5pi/15 - 3pi/15) = cos (2 pi/15) Likewise, you can switch your A and B since A and B are commutative so you cos (-2pi/15) Both give you the same value. Remember the unit circle. cos is the x coordinate and sin is the y coordinate. 2pi/15 is in the first quadrant and -2pi/15 is in the 4th quadrant.

OpenStudy (anonymous):

So that is the whole answer? Thanks for replying man.

OpenStudy (anonymous):

Yes, the angle is 2pi/15 or -2pi/15. It might also be presented in a different form other than -2pi/15 since -2pi/15 is just 2pi - 2pi/15. (the unit circle repeats every 2 pi.)

OpenStudy (anonymous):

in radians of course

OpenStudy (anonymous):

Thank you so much man. I really appreciate it.